OpenStudy (lilshane):

Which three ordered pairs complete the table to give solutions of the equation y = x – 7? A. (–3, –10), (0, –7), and (3, 4) B. (–3, –10), (0, –7), and (3, –4) C. (–3, –4), (0, –7), and (3, –4) D. (–3, 4), (0, –7), and (3, –4)

OpenStudy (lilshane):

@inowalst @EclipsedStar @eta

OpenStudy (lilshane):

@just_one_last_goodbye @JFraser @AutumnRoseT @sissyedgar @leonardo0430 @Mehek14 @prettygirl876

Mehek (mehek14):

you have to plug in the ordered pairs

OpenStudy (lilshane):

i know

Mehek (mehek14):

Lets start with a

OpenStudy (just_one_last_goodbye):

okay, which one do you think it is?

OpenStudy (lilshane):

everyone keeps saying that but i know know how

Mehek (mehek14):

the first ordered pair is (-3,-10) -3 is x -10 is y -10 = -3 - 7 is that true^?

OpenStudy (just_one_last_goodbye):

(x,y) so plug the numbers (-3,-10) where x is -3 and y is -10

OpenStudy (lilshane):

no?

Mehek (mehek14):

what is -3-7?

OpenStudy (lilshane):

wait that is true sorry

Mehek (mehek14):

ok so so far it could be A but we have to check the rest of the ordered pairs

OpenStudy (lilshane):

ok

Mehek (mehek14):

next ordered pair (0,-7) 0=x -7=y -7=0-7 is that true?

OpenStudy (lilshane):

n o

Mehek (mehek14):

0-7?

OpenStudy (lilshane):

omg i keep getting confused lol

Mehek (mehek14):

so is it true?

OpenStudy (lilshane):

yes

Mehek (mehek14):

ok now last ordered pair

Mehek (mehek14):

(3,4) 3=x 4=y 4=3-7 true or not?

OpenStudy (lilshane):

false

Mehek (mehek14):

ok so now move on to B

OpenStudy (lilshane):

the first two are true

Mehek (mehek14):

the first 2 ordered pairs are the same as in A so look at the last one

Mehek (mehek14):

(3,-4) -4=3-7

Mehek (mehek14):

is that true

OpenStudy (lilshane):

yes! so its B?

Mehek (mehek14):

yup

OpenStudy (lilshane):

thanks you

Mehek (mehek14):

yw

OpenStudy (lilshane):

mind staying here and help me go over the next question?

OpenStudy (lilshane):

?